OpenStudy (anonymous):

An equilateral triangle has an altitude length of 36 feet. Determine the length of a side of the triangle.

OpenStudy (anonymous):

41.57?

OpenStudy (lgbasallote):

\[\frac{x^2}{4} + 36^2 = x^2\] \[x^2 + 5184 = 4x^2\] \[5184 = 3x^2\] \[1728 = x^2\] \[x = 41.57\]